OpenStudy (anonymous):

What is the diameter of a circle with area equal to one-hundred sixty nine pi meters squared?

OpenStudy (lgbasallote):

\[\large \sqrt{\frac{4A}{\pi}} = d\] so sub in your values

OpenStudy (lgbasallote):

do you need any further help?

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

if you wanna know how i got that equation... \[A = \pi (\frac{d}{2})^2\] \[A = \pi \frac{d^2}{4}\] \[\frac{4A}{\pi} = d^2\] \[\sqrt{\frac{4A}{\pi}} = d\]

OpenStudy (lgbasallote):

so to sub in...

OpenStudy (lgbasallote):

\[\large \sqrt{\frac{4(169 \cancel{\pi})}{\cancel{\pi}}} = d\] \[d = \sqrt{4 \times 169}\] \[d = 2 \times 13\] \[d = 26\]

OpenStudy (anonymous):

thanks for explaining it